# Current-Enabled Optical Conductivity of Collective Modes in Unconventional Superconductors

### v2

@@ V Discussion
We first applied these to toy models with $s$-wave and $d$-wave pairing interactions. In these examples, various collective modes were found to become optically active in the presence of an injected supercurrent. A one-band model exhibits a peak associated with a BS mode in the $s$-wave phase, which softens at the transition and turns into an MSBS mode of the $s-id$-fluctuations in the $s+id$-phase. This peak stays visible throughout the entire phase, although the intensity changes depending on the mode frequency. This collective mode also strongly suppresses the known peak [30] at the quasiparticle gap, which happened for all the models studied in this paper, including the Rashba system (where no supercurrent was needed to induce a non-trivial optical response). The presence of a BS mode always seems to suppress the quasiparticle contribution in the optical conductivity, an effect which becomes stronger as the BS mode lies further below the gap.
@@
We then studied this situation in a simple two-band model. Neglecting the interband pairing, we found that considering interband pair scattering was sufficient to generate qualitative changes in the collective-mode spectrum, which now contained a Leggett mode and two BS modes: One which is in phase, one which is out of phase between the two bands. We found that the Leggett mode and in-phase BS mode always generate a peak in the optical conductivity, while the out-of-phase mode only appears when the gaps do not coinside between the two bands.
@@
This could be applied to the study of iron-based superconductors (FeSC), which are thought to feature competing $s$-wave and $d$-wave pairing interactions [3]. The BS modes have been observed in such materials using Raman spectroscopy [48, 49]. The current-enabled linear optical conductivity will provide a new way of studying and better understanding the pairing mechanism. However, it should be mentioned that the optical conductivity does not provide information about the symmetry of the mode. This is in contrast to the Raman spectroscopy, where the signature of a collective mode can be separated in different channels.
Lastly, we considered a model for $p$-wave superconductors with the interband pairing and the Rashba-type SOC, again with several pairing channels. Here, we found that the model exhibits peaks due to BS modes in the linear optical response even without a supercurrent. In a chiral $p$-wave phase, these peaks are also seen in the ac Hall conductivity, suggesting possible experimental observation via the polar Kerr effect [46, 47].
@@
Note that in this model, due to the previously calculated Lifshitz invariants [35], finite-momentum Cooper pairing could be realized without the need for considering a supercurrent. Regardless of the physical origin, we also found that in the chiral $p$-wave phase, finite-momentum Cooper pairing allows for a peak in the longitudinal optical conductivity due to relative phase oscillations of the two order parameter components [37]. This is not associated with the Rashba system specifically, and remains for a chiral $p$-wave order parameter without any spin-orbit coupling. Therefore, this provides a new experimental probe for chiral $p$-wave superconductors, or chiral $d$-wave superconductors which feature the same mode [37]. Sr2RuO4, for example, has often been theorized as an example for such states, but this has been difficult to verify experimentally [50]. Optical measurements in the presence of an injected supercurrent therefore present a promising avenue for exploring the ground state of this and related materials. All of these suggests a promising way of exploring the properties of unconventional superconductors, in particular those suspected to feature sub-dominant pairing channels or multiple superconducting bands. The calculated formulas are fairly general, and can easily be applied to any band structures and pairing interactions, allowing comparison with experimental data.
@@
There also remain several open questions for further theoretical investigation. Firstly, the mechanism behind the suppression of the quasiparticle peak by the BS modes should be investigated. Secondly, disorder in superconductors may have significant effects on the optical conductivity, such as introducing Mattis-Bardeen terms above the superconducting gap [51]. It is open to understand in what way the presence of disorders would modify our results.
@@ Appendix C Quench dynamics in the one-band model with $s$- and $d$-wave pairing interactions
We adopt two methods for the quench simulation: The first one is a “$\Delta$-quench”, in which a small fluctuation from equilibrium is introduced in the initial gap function, and then we let the system evolve from the corresponding initial state. The second method is a “$V$-quench”, in which the interaction parameters $V_{s}$ and $V_{d}$ are suddenly changed. We use both methods here, since it seems impossible to induce the BS mode in the pure $s$-wave phase by the $V$-quench, and so is it to discuss the damping behavior based on the linearized analysis for the $\Delta$-quench. We confirm that the oscillation behavior of the Higgs and MSBS modes are consistent between these two quench methods.
@@ C.1 Equilibrium phase diagram
We solve Eq. (97) under the assumption that the $s+d$ solution is unstable [43]. There are three types of non-trivial solutions: pure-$s$, pure-$d$, and $s+id$. Figure 22 shows the equilibrium phase diagram. We will later show the results of the quench dynamics for the set of parameters, $(V_{s},V_{d})=(-5,-6)$, $(-5,-10)$ (marked by red and blue crosses in Fig. 22), which correspond to the pure-$s$ and $s+id$ phases, respectively.
Figure 22: Equilibrium phase diagram of the one-band model with the $s$- and $d$-wave pairing interactions studied in Sec. III.1. There are three superconducting phases (pure-$s$, pure-$d$, and $s+id$) as well as the normal state. The red and blue markers represent $(V_{s},V_{d})=(-5,-6)$ and $(-5,-10)$, used in the simulation later. The parameters are $\mu=0$, $t=1$, and $T=0.05$.
@@ C.3.1 Pure $s$-wave phase
is obtained. In the range of $0<\omega<2\Delta_{s}^{\textrm{eq}}$, the derivative $\frac{dD}{d\omega}$ is positive. It is not difficult to show that $D(0)<-1/V_{d}$ and $\lim_{\omega\to 2\Delta_{s}^{\textrm{eq}}}D(\omega)=+\infty$. Therefore, Eq. (115) has a unique solution in $0<\omega<2\Delta_{s}^{\textrm{eq}}$. Here, $D(\omega)$ crosses $-1/V_{d}$ with a finite gradient. This frequency corresponds to that of the BS mode.
@@ C.4 Undamped behavior of the collective modes
×$\quad\times\begin{pmatrix}\frac{\Delta_{s}^{\textrm{eq}}}{\mp i\omega_{0}}\frac{\delta V_{s}}{V_{s}}\\ \frac{\Delta_{d}^{\textrm{eq}}}{\mp i\omega_{0}}\frac{\delta V_{d}}{V_{d}}\end{pmatrix}.$ (127)
@@ Appendix D The Rashba gap in the spin basis
$=\frac{1}{2}\begin{pmatrix}t(\widetilde{\Delta}_{++}+\widetilde{\Delta}_{-+}-\widetilde{\Delta}_{+-}-\widetilde{\Delta}_{--})&-\widetilde{\Delta}_{++}+\widetilde{\Delta}_{-+}+\widetilde{\Delta}_{+-}-\widetilde{\Delta}_{--}\\ +\widetilde{\Delta}_{++}-\widetilde{\Delta}_{-+}-\widetilde{\Delta}_{+-}+\widetilde{\Delta}_{--})&t^{*}(-\widetilde{\Delta}_{++}+\widetilde{\Delta}_{-+}-\widetilde{\Delta}_{+-}+\widetilde{\Delta}_{--})\end{pmatrix}.$ (133)
@@
$\begin{split}\widehat{\Delta}^{\uparrow\downarrow}&=\psi(\bm{k})i\sigma_{y}+i(\bm{d(k)\cdot\sigma})\sigma_{y}\\ &=\begin{pmatrix}-d_{x}(\bm{k})+id_{y}(\bm{k})&d_{z}(\bm{k})+\psi(\bm{k})\\ d_{z}(\bm{k})-\psi(\bm{k})&d_{x}(\bm{k})+id_{y}(\bm{k})\end{pmatrix},\end{split}$ (134)

### v3

@@ V Discussion
We first applied these to toy models with $s$-wave and $d$-wave pairing interactions. In these examples, various collective modes were found to become optically active in the presence of an injected supercurrent. A one-band model exhibits a peak associated with a BS mode in the $s$-wave phase, which softens at the transition and turns into an MSBS mode of the sids-id-fluctuations in the $s+id$-phase. This peak stays visible throughout the entire phase, although the intensity changes depending on the mode frequency. This collective mode also strongly suppresses the known peak [31] at the quasiparticle gap, which happened for all the models studied in this paper, including the Rashba system (where no supercurrent was needed to induce a non-trivial optical response). The presence of a BS mode always seems to suppress the quasiparticle contribution in the optical conductivity, an effect which becomes stronger as the BS mode lies further below the gap.
@@
We then studied this situation in a simple two-band model. Neglecting the possibility of interband cooper pairs, we found that interband scattering of intraband cooper pairs was sufficient to generate qualitative changes in the collective-mode spectrum, which now contained a Leggett mode and two BS modes: One which is in phase, one which is out of phase between the two bands. We found that the Leggett mode and in-phase BS mode always generate a peak in the optical conductivity, while the out-of-phase mode only appears when the gaps do not coinside between the two bands.
@@
This could be applied to the study of iron-based superconductors (FeSC), which are thought to feature competing $s$-wave and $d$-wave pairing interactions [3]. The BS modes have been observed in such materials using Raman spectroscopy [49, 50]. The current-enabled linear optical conductivity will provide a new way of studying and better understanding the pairing mechanism. However, it should be mentioned that the optical conductivity does not provide information about the symmetry of the mode. This is in contrast to the Raman spectroscopy, where the signature of a collective mode can be separated in different channels. Furthermore, iron-based superconductors are also thought to feature a greater variety of $s$-wave ground states such as $_{\pm}, s++s_{++}$ or TRS-breaking combinations like $s_{\pm}+is_{++}$ can arise [3]. Whether such states also feature optically active collective modes in the presence of a supercurrent remains an open question. Lastly, we considered a model for $p$-wave superconductors with the interband pairing and the Rashba-type SOC, again with several pairing channels. Here, we found that the model exhibits peaks due to BS modes in the linear optical response even without a supercurrent. In a chiral $p$-wave phase, these peaks are also seen in the ac Hall conductivity, suggesting possible experimental observation via the polar Kerr effect [47, 48].
@@
Note that in this model, due to the previously calculated Lifshitz invariants [36], finite-momentum Cooper pairing could be realized without the need for considering a supercurrent. Regardless of the physical origin, we also found that in the chiral $p$-wave phase, finite-momentum Cooper pairing allows for a peak in the longitudinal optical conductivity due to relative phase oscillations of the two order parameter components [38]. This is not associated with the Rashba system specifically, and remains for a chiral $p$-wave order parameter without any spin-orbit coupling. Therefore, the presence (or absence) of such a current-enabled peak provides a new experimental probe for confirming (or ruling out) the chiral $p$-wave symmetry of an order parameter. However, as mentioned before, the peaks contain no information about the symmetry of collective modes, so this would have to be combined with other experimental probes. Sr2RuO4, for example, has often been theorized as an example for such states, but this has been difficult to verify experimentally [51]. Optical measurements in the presence of an injected supercurrent therefore present a promising avenue for exploring the ground state of this and related materials. All of these suggests a promising way of exploring the properties of unconventional superconductors, in particular those suspected to feature sub-dominant pairing channels or multiple superconducting bands. The calculated formulas are fairly general, and can easily be applied to any band structures and pairing interactions, allowing comparison with experimental data.
@@
There also remain several open questions for further theoretical investigation. Firstly, the mechanism behind the suppression of the quasiparticle peak by the BS modes should be investigated. Secondly, disorder in superconductors may have significant effects on the optical conductivity, such as introducing Mattis-Bardeen terms above the superconducting gap [52].
It has been shown that contributions from the Higgs mode to the optical conductivity exist in the presence of disorder, so we would expect that collective modes below the gap will also survive after introducing impurities. Further research is needed to verify this and to quantify the effect of disorder on our results.
@@ Appendix C Quench dynamics in the one-band model with $s$- and $d$-wave pairing interactions
We adopt two methods for the quench simulation: The first one is a $\Delta$-quench”, in which a small fluctuation from equilibrium is introduced in the initial gap function, and then we let the system evolve from the corresponding initial state. The second method is a “$V$-quench”, in which the interaction parameters $V_{s}$ and $V_{d}$ are suddenly changed. We use both methods here, since it seems impossible to induce the BS mode in the pure $s$-wave phase by the $V$-quench, and so is it to discuss the damping behavior based on the linearized analysis for the $\Delta$-quench. We confirm that the oscillation behavior of the Higgs and MSBS modes are consistent between these two quench methods.
@@ C.1 Equilibrium phase diagram
We solve Eq. (101) under the assumption that the $s+d$ solution is unstable [44]. There are three types of non-trivial solutions: pure-$s$, pure-$d$, and $s+id$. Figure 22 shows the equilibrium phase diagram. We will later show the results of the quench dynamics for the set of parameters, $(V_{s},V_{d})=(-5,-6)$, (5,10)(-5,-10) (marked by red and blue crosses in Fig. 22), which correspond to the pure-$s$ and $s+id$ phases, respectively.
Figure 22: Equilibrium phase diagram of the one-band model with the $s$- and $d$-wave pairing interactions studied in Sec. III.1. There are three superconducting phases (pure-$s$, pure-$d$, and $s+id$) as well as the normal state. The red and blue markers represent $(V_{s},V_{d})=(-5,-6)$ and (5,10)(-5,-10), used in the simulation later. The parameters are $\mu=0$, $t=1$, and $T=0.05$.
@@ C.3.1 Pure $s$-wave phase
is obtained. In the range of $0<\omega<2\Delta_{s}^{\textrm{eq}}$, the derivative $\frac{dD}{d\omega}$ is positive. It is not difficult to show that $D(0)<-1/V_{d}$ and $\lim_{\omega\to 2\Delta_{s}^{\textrm{eq}}}D(\omega)=+\infty$. Therefore, Eq. (119) has a unique solution in $0<\omega<2\Delta_{s}^{\textrm{eq}}$. Here, $D(\omega)$ crosses $1/V_{d}$ with a finite gradient. This frequency corresponds to that of the BS mode.
@@ C.4 Undamped behavior of the collective modes
$\quad\times\begin{pmatrix}\frac{\Delta_{s}^{\textrm{eq}}}{\mp i\omega_{0}}\frac{\delta V_{s}}{V_{s}}\\ \frac{\Delta_{d}^{\textrm{eq}}}{\mp i\omega_{0}}\frac{\delta V_{d}}{V_{d}}\end{pmatrix}$. (131)
@@ Appendix D The Rashba gap in the spin basis
[…]
@@
[…]
@@
Appendix E Free energy calculations
In cases where several ground states fulfill the gap equation (see Appendix A), a free energy calculation is necessary to find the state which will be realized in the system. This can be done using the usual formula for the free energy of non-interacting fermions[39]
$\mathcal{F}$ $=-\frac{1}{\beta}\sum_{\bm{k}j}\ln\left(1+e^{-\beta E^{j}_{\bm{k}}}\right),$ (140)
where $\beta$ is the inverse temperature and $j$ are any quantum numbers that need to be summed over in addition to momentum $\bm{k}$. At zero temperature this simplifies to
$\mathcal{F}$ $=\sum_{\bm{k},j:E^{j}_{\bm{k}}<0}E^{j}_{\bm{k}}$ (141)
which is simply the total energy of all occupied states. For the systems with $s$- and $d$-wave interactions, the main competition lies in between the pure $s$-wave ground state and the mixed-symmetry $s+id$-wave ground state. As the interaction in the $d$-wave channel is increased, the pure $s$-wave state always remains a solution of the gap equation, so the free energies of the two states should be compared. For the example of the multiband system with asymmetrical bands, this comparison is shown in Fig. 26.
Figure 26: Free energy of the multiband system with asymmetrical bands (Sec. III.2.2), which is analogous to the symmetrical two-band case (Sec. III.2.1) and the one-band case (Sec III.1). Parameters are again $\mu=-0.2$, $V_{s}=-1$, $t=1$, and $\bm{q}=0.001\bm{\hat{e}}_{x}$.
This graph shows how the free energies of the mixed and pure symmetry state initially coincide, when the ”mixed” symmetry state has no $d$ -wave component, and they are identical. As soon as the states become different, one can observe that the free energy of the mixed state always lies below that of the pure $s$-wave solution to the gap equation. The plot is shown for the most complicated case of two asymmetrical bands, but is identical for the other cases which were discussed in the article.
For the Rashba system (see Sec. IV.3), the transition we considered was between an $A_{1}$ ground state, corresponding to an in-plane $p$-wave $\bm{d}$-vector, and an out-of-plane, chiral $p_{x}+ip_{y}$-wave $\bm{d}$-vector. Additionally, a real out-of-plane $p_{x}+p_{y}$ state would have solved the relevant gap equation. The comparison of the relevant free energies is shown in Fig. 27.
Figure 27: Free energies of possible ground states belonging to the $E$- or $A_{1}$-irrep of the Rashba system (Sec. IV), for increasing $V_{E}$. Parameters are $\mu=-3$, $t=1$, $\alpha=0.1$, $\bm{q}=0$, and $\omega_{c}=1$ and $V_{A_{1}}=1$.
This shows how the free energy of the chiral ground state in the $E$-irrep crosses that of the $A_{1}$ ground state at the transition. It can also be seen how the free energy of the real $p_{x}+p_{y}$ state in the $E$ irrep eventually falls below that of the $A_{1}$ state, but remains above that of the chiral state. This was used to confirm the phase diagram in Fig. 17a.